Please solve this-


 


In (a=0, b=1)    {  [log(1+x)]  /  (1+x^2)   }  dx

Grade:

1 Answers

SREEDHAR P
11 years ago

put x=tana then dx=sec^2ada , integral 0 to pi/4  log(1+tana)sec^2a/1+tan^2a

 = log(1+tana) on applying integral a-b f(x) = integral a-b f(a+b-x) we get required definite integral as( pi/8)log2